# Kink Finder at Belle II

### v1

@@ 1 Introduction
The Belle II experiment is conducted at the SuperKEKB accelerator in Tsukuba, Japan. SuperKEKB is a high-luminosity asymmetric energy $e^{+}e^{-}$ collider with a 4 GeV positron beam and a 7 GeV electron beam, providing an optimal center-of-mass energy of around 10.58 GeV. The Belle II experiment enables precision studies of flavor physics and rare decays and direct searches for physics beyond the Standard Model (e.g., forbidden decays, dark sector particles), including through the application of novel methods. These goals require the development of high-quality hardware and software.
@@
The reconstruction of a so-called kink – the signature of a particle that decays or scatters in-flight, leading to a sudden and significant change in the direction of motion – is crucial to improve the performance of the experiment. Furthermore, it increases the scope of the experiment, enabling the measurement of hitherto unmeasured physics processes. Processes that result in kinks include $K^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}$ [^1] Charge conjugation is implied throughout the paper unless otherwise indicated., $K^{-}\rightarrow\pi^{-}\pi^{0}$, $K^{-}\rightarrow\pi^{-}\pi^{0}\pi^{0}$, $K^{-}\rightarrow\pi^{-}\pi^{+}\pi^{-}$, $K^{-}\rightarrow\pi^{0}\mu^{-}\bar{\nu}_{\mu}$, $K^{-}\rightarrow\pi^{0}e^{-}\bar{\nu}_{e}$, $\pi^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}$, and $\mu^{-}\rightarrow e^{-}\bar{\nu}_{e}\nu_{\mu}$, but also scattering in the detector material. Kink reconstruction can reduce the misidentification (fake) rate in the particle identification (PID) of the experiment. For example, $K^{-}\rightarrow\pi^{-}X$ in-flight decay can be reconstructed as one track, where the hits from the daughter pions in the outer detector systems will dominate the reconstruction, hence misidentifying the mother kaon as a pion. In this paper, we refer to the primary track of the kink as mother and the secondary track of the kink as daughter. Furthermore, two-body kaon and pion decays can be used to improve the $dE/dx$ calibration, since the mother particle can be identified by a peak in the momentum distribution of the daughter particle in the mother’s rest frame.
@@
A well-known feature of the track-finding software is that it accidentally reconstructs the trajectory of one single particle as two separate tracks, creating a pair of duplicates known as clone tracks. The point where these tracks intersect can mimic a kink vertex. Identifying and removing these clones improves tracking performance by reducing the double counting of a single particle.
@@
Finally, kinks are interesting for physics, e.g., muon in-flight decay reconstruction provides information about the muon polarization, which is important for the Michel parameters measurement in $\tau^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}\nu_{\tau}$ decays [Bodrov:2022mbd, Belle:2023udc, Belle:2023dyc].
@@
To our best knowledge, not many experiments have implemented an algorithm for searching for kinks; the most well-documented algorithm is provided by the NOMAD Collaboration [NOMAD:1997pcg] in Ref. [Astier:1999rs]. Unfortunately, the differences between the Belle II and the NOMAD experimental setups, in combination with the specific structure of the Belle II tracking software, including the genfit2 track fitting package, does not allow for direct implementation of this algorithm. Thus, Belle II needs its own, dedicated algorithm.
@@ 2 Belle II experiment
The Belle II experiment is located at the SuperKEKB asymmetric-energy $e^{+}e^{-}$ collider [Akai:2018mbz]. The Belle II detector consists of subsystems arranged cylindrically around the interaction point (IP) [Abe:2010gxa, Kou:2018nap]. Belle II uses right-handed orthogonal coordinates in which the $z$-axis is approximately collinear with the electron beam, the $x$-axis points horizontally away from the center of the accelerator rings, and the $y$-axis points vertically upward, perpendicular to the accelerator plane. In practice, cylindrical coordinates are used with the radial distance $r=\sqrt{x^{2}+y^{2}}$ and the azimuthal angle $\varphi$ in the $xy$-plane, measured from the $x$-axis. Hits used for reconstruction of charged-particle trajectories (tracks) are measured by a two-layer silicon-pixel detector (PXD) surrounded by a four-layer double-sided silicon-strip vertex detector (SVD) and a central 56-layer drift chamber (CDC). The latter two detectors also measure the energy loss from ionization. A quartz-based Cherenkov detector (TOP) identifies charged hadrons in the central region by measuring the time-of-propagation of the Cherenkov light emitted when the particles cross the quartz bars. An aerogel-based ring-imaging Cherenkov counter (ARICH) identifies charged hadrons in the forward region. An electromagnetic calorimeter (ECL) made of CsI(Tl) crystals measures photon and electron energies and directions. The above subdetectors are immersed in a 1.5 T axial magnetic field provided by a superconducting solenoid. A subdetector dedicated to identifying muons and $K_{L}^{0}$ mesons (KLM) is installed outside of the solenoid.
@@
The most important subdetector for the kink reconstruction is the CDC, since part of the mother track and the full daughter track are reconstructed based on CDC hits. The CDC inner (outer) radius is 17 (110) cm. It consists of 56 layers arranged in nine superlayers. The superlayers with wires aligned parallel to the solenoid magnetic field are called axial (A). To provide three-dimensional (3D) tracking, the stereo superlayers are used. They contain wires skewed by an angle between 66.8 and 74.1 mrad in the positive direction and 58-58 to 78.6-78.6 mrad in the negative direction, designated U and V, respectively. The resulting sequence of superlayers is A, U, A, V, A, U, A, V, A.
@@
For simulation, reconstruction, calibration, and physics analysis within Belle II, a central, modular C++/Python framework, called Belle II Analysis Software Framework (basf2), is used [Kuhr:2018lps, basf2-zenodo]. For the track reconstruction, Belle II employs a multi-step scheme implemented within basf2. The process begins with a CDC-only track-finding based on the Legendre algorithm [Alexopoulos:2008zza], followed by attachment of SVD clusters with a combinatorial Kalman filter (CKF) [Mankel:1998uy, Mankel:2004yv, CMS:2014pgm, ATLAS:2017kyn]. The remaining SVD hits are used to reconstruct tracks with a standalone SVD track-finder using Sector Map and cellular automaton algorithms [Gardner:1970, Glazov:1993ur, Abt:2002he, Funke:2014dga]. The results are combined and extrapolated to the PXD with a second CKF. Prior to the extrapolation between different detectors and for the final combination, the tracks are fitted using the deterministic annealing filter (DAF) provided by the genfit2 package [Bilka:2019ang, jojosito_2023_10358638]. More detailed information about the Belle II track-finding algorithm is provided in Ref. [Bertacchi:2020eez].
@@
The benchmarking of the Kink Finder is carried out using $\tau^{+}\tau^{-}$ events, generated with KKMC [Jadach:1999vf] and TAUOLA [Jadach:1990mz] packages, and $B\bar{B}$ events, generated with PYTHIA8 [Sjostrand:2014zea] and EvtGen [Lange:2001uf] packages. The full detector geometry, interactions of final state particles with the detector material, and the detector response are simulated using the GEANT4 package [Agostinelli:2002hh].
@@ 3 Kinks at Belle II
The latter case is relevant to the PID fake rate, as mentioned in Sec. 1. In around 10% of in-flight decay events, the daughter track cannot be reconstructed since it leaves too few hits inside the CDC. Although some of the kink processes produce more than one charged daughter particle, it is usually possible to reconstruct and combine only one of them with the mother track, e.g., it is quite rare that more than one daughter pion is reconstructed from the $K^{-}\rightarrow\pi^{-}\pi^{+}\pi^{-}$ decay. Examples of MC simulated events with kinks inside the Belle II CDC are shown in Fig. 1. The blue and red hits show the reconstructed mother and daughter tracks, respectively, while yellow hits represent all reconstructed hits inside the CDC, including the background.
@@
There are around 40% of kaon and 70% of pion events containing in-flight decays, in which there is one reconstructed track arising from a combination of mother and daughter hits. Often, the hits attributed to either the mother or the daughter dominate in such combined track, allowing for its relation to the dominant particle at the MC level. Therefore, some of the combined tracks are included in the calculation of the reconstruction efficiencies above. Around 40% of the combined tracks can neither be attributed to the mother nor to the daughter. This is because the ratio of the hit contributions is almost 50%/50%, whereas the Belle II definition requires at least 66% of hits from a single particle to create a relation with it [Bertacchi:2020eez].
@@ 4 Kink Finder algorithm
4.1 Selection of the kink candidates
To optimize the execution time of the algorithm, we perform a preselection of tracks that are candidates to form a kink. First, we describe the selection for the case with both mother and daughter tracks reconstructed individually. The mother track candidate should start in the vicinity of the IP and, if it has CDC hits, end inside the CDC. The daughter track candidate should start (or end, if the assigned charge of the track is wrong) inside the CDC and its point of closest approach to the IP should be at a minimum distance.
@@ 4.1 Selection of the kink candidates
    The starting point of the daughter track is close to the end-point of the mother track in 3D [see Fig. 1(a)].
@@
    The end-point of the daughter track is close to the end-point of the mother track in 3D [see Fig. 1(b)]. This configuration covers events with wrong charge assignment of the daughter track.
@@
    The extrapolated daughter track is close to the end-point of the mother track in 3D [see Fig. 1(c)]. This configuration covers events in which the daughter track intermediate hits are lost.
@@
  1. 1.
    The p-value of the track fit must be lower than the threshold, which is set several orders of magnitude larger than the loose threshold used in general track fitting. The value used in the Kink Finder is optimized for the experimental conditions to suppress the splitting of one true track into two false tracks.
  2. 2.
  3. 3.
After tracks are selected as kink candidates, they are processed to reconstruct the kink vertex, determine the kinematics of the daughter and mother tracks at the kink vertex, filter out fake candidates, and store the selected kinks in a specially designed kink object used further at the analysis level.
4.2 Reconstructed track pairs
In the case of two separately reconstructed tracks, several assisting algorithms are applied to improve the kink parameters resolution. One example is hit reassignment between tracks and daughter track refitting with better initial values (seeds).
We start with track pairs selected by the first three configurations (described in Sec. 4.1), all of which have good $z$-coordinate resolution for the daughter track. For these pairs, the processing begins with a vertex fit to find the location of the kink. The fitting procedure is described in Sec. 4.2.1. In the case of poor $z$-coordinate resolution of the daughter track, the track is refitted using the time and position information from the endpoint of the mother track. If the track charges are correctly reconstructed, and the start point and end point of the tracks are close to each other (normal ordering), the poor $z$-coordinate resolution of the daughter track may be due to wrong assignment of the first stereo superlayer in the CDC. We then try to refit the daughter track without these hits. Using the new fit result for the daughter, the vertex fit is performed.
@@ 4.2 Reconstructed track pairs
While the DAF fitting algorithm from the genfit2 package is used as the default track fitter at Belle II, we use the Reference Kalman Filter from the same package [Bilka:2019ang, jojosito_2023_10358638] for the fit of the flipped tracks and the refit of the daughter tracks with missing superlayer hits, as it shows a better performance.
@@
      The p-value of the fitted combined track is greater than the initial mother track’s p-value.
@@ 4.2.1 Vertex fit
The fit of the vertex position is performed using a kinematic fitter algorithm implemented in the Belle II Software. It is based on the least squares method and described in Ref. [Tanaka:2001ae], Appendix D. The genfit2 package also provides an algorithm for the vertex fitting, GFRave [Bilka:2019ang, jojosito_2023_10358638, Waltenberger:2011zz], which is applied at Belle II for the V0 ($K^{0}_{S}$, $\Lambda$, and externally converted $\gamma$) vertex reconstruction. However, we do not use it in the Kink Finder since it does not have an option for setting an initial value. Without such an option, the vertex fit fails for a noticeable amount of kinks.
@@ 4.3 Combined tracks
The combined track candidates are split using the algorithm that creates a pair of mother and daughter tracks using the hits of the initial track.
@@ 5 Performance
The Kink Finder performance for in-flight decays is studied using MC samples of the $\tau^{-}\rightarrow K^{-}\nu_{\tau}$ and $\tau^{-}\rightarrow\pi^{-}\nu_{\tau}$ events with a total number of $8\times 10^{4}$ kaon and $13\times 10^{4}$ pion in-flight decays inside the CDC. The performance for general cases of kinks is studied using MC samples of generic $\tau$-pair and $B\bar{B}$ decays with $10^{5}$ events in each sample. In both studies, we account for beam-induced background, which consists of spurious detector hits and particles produced in secondary interactions with the detector and support material outside the IP region [Natochii:2022vcs]. The level of this background strongly depends on the SuperKEKB luminosity and its conditions. In our studies, we use a configuration that is considered typical for current Belle II operations. Although the Kink Finder has been included in the Belle II tracking software, it has not been applied for data processing yet, and therefore, there is no performance study using data available at the time of writing.

### v2

@@ 1 Introduction
The Belle II experiment is conducted at the SuperKEKB accelerator in Tsukuba, Japan. SuperKEKB is a high-luminosity asymmetric energy $e^{+}e^{-}$ collider with a 4 GeV positron beam and a 7 GeV electron beam, providing a center-of-mass energy of around 10.58 GeV. The Belle II experiment enables precision studies of flavor physics and rare decays and direct searches for physics beyond the Standard Model (e.g., forbidden decays, dark sector particles), including through the application of novel methods. These goals require the development of high-quality hardware and software.
@@
The reconstruction of a so-called kink – the signature of a particle that decays or scatters in-flight, leading to a sudden and significant change in the direction of motion – is one of the steps to improve the performance of the experiment. Furthermore, it increases the scope of the experiment, enabling the measurement of hitherto unmeasured physics processes. Processes that result in kinks include $K^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}$ [^1] Charge conjugation is implied throughout the paper unless otherwise indicated., $K^{-}\rightarrow\pi^{-}\pi^{0}$, $K^{-}\rightarrow\pi^{-}\pi^{0}\pi^{0}$, $K^{-}\rightarrow\pi^{-}\pi^{+}\pi^{-}$, $K^{-}\rightarrow\pi^{0}\mu^{-}\bar{\nu}_{\mu}$, $K^{-}\rightarrow\pi^{0}e^{-}\bar{\nu}_{e}$, $\pi^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}$, and $\mu^{-}\rightarrow e^{-}\bar{\nu}_{e}\nu_{\mu}$, but also scattering in the detector material. Kink reconstruction can reduce the misidentification (fake) rate in the particle identification (PID) of the experiment. For example, $K^{-}\rightarrow\pi^{-}X$ in-flight decay can be reconstructed as one track, where the hits from the daughter pions in the outer detector systems will dominate the reconstruction, hence misidentifying the mother kaon as a pion. In this paper, we refer to the primary track of the kink as mother and the secondary track of the kink as daughter. Furthermore, two-body kaon and pion decays can be used to improve the $dE/dx$ calibration, since the mother particle can be identified by a peak in the momentum distribution of the daughter particle in the mother’s rest frame.
@@
A well-known feature of the track-finding software is that it may accidentally reconstruct the trajectory of one single particle as two separate tracks, creating a pair of duplicates known as clone tracks. The point where these tracks intersect can mimic a kink vertex. Identifying and removing these clones improves tracking performance by reducing the double counting of a single particle.
@@
Finally, kinks are interesting for physics, e.g., muon in-flight decay reconstruction provides information about the muon polarization, which is important for the Michel parameters measurement in $\tau^{-}\rightarrow\mu^{-}\bar{\nu}_{\mu}\nu_{\tau}$ decays [16, 14, 15].
@@
To our best knowledge, not many experiments have implemented an algorithm for searching for kinks; the most well-documented algorithm is provided by the NOMAD Collaboration [8] in Ref. [10]. Unfortunately, the differences between the Belle II and the NOMAD experimental setups, in combination with the specific structure of the Belle II tracking software, including the genfit2 track fitting package, does not allow for direct implementation of this algorithm. Thus, Belle II needs its own, dedicated algorithm.
@@ 2 Belle II experiment
The Belle II experiment is located at the SuperKEKB asymmetric-energy $e^{+}e^{-}$ collider [6]. The Belle II detector consists of subsystems arranged cylindrically around the interaction point (IP) [2, 9]. Belle II uses right-handed orthogonal coordinates in which the $z$-axis is approximately collinear with the electron beam, the $x$-axis points horizontally away from the center of the accelerator rings, and the $y$-axis points vertically upward, perpendicular to the accelerator plane. In practice, cylindrical coordinates are used with the radial distance $r=\sqrt{x^{2}+y^{2}}$ and the azimuthal angle $\varphi$ in the $xy$-plane, measured from the $x$-axis. Hits used for reconstruction of charged-particle trajectories (tracks) are measured by a two-layer silicon-pixel detector (PXD) surrounded by a four-layer double-sided silicon-strip vertex detector (SVD) and a central 56-layer drift chamber (CDC). The latter two detectors also measure the energy loss from ionization. A quartz-based Cherenkov detector (TOP) identifies charged hadrons in the central region by measuring the time-of-propagation of the Cherenkov light emitted when the particles cross the quartz bars. An aerogel-based ring-imaging Cherenkov counter (ARICH) identifies charged hadrons in the forward region. An electromagnetic calorimeter (ECL) made of CsI(Tl) crystals measures photon and electron energies and directions. The above subdetectors are immersed in a 1.5 T axial magnetic field provided by a superconducting solenoid. A subdetector dedicated to identifying muons and $K_{L}^{0}$ mesons (KLM) is installed outside of the solenoid.
@@
The Kink Finder aims to reconstruct kinks inside the CDC. Therefore, a part of the mother track and the full daughter track are reconstructed based on CDC hits. The CDC inner (outer) radius is 17 (110) cm. It consists of 56 layers arranged in nine superlayers. The superlayers with wires aligned parallel to the solenoid magnetic field are called axial (A). To provide three-dimensional (3D) tracking, the stereo superlayers are used. They contain wires skewed by an angle between 66.8 and 74.1 mrad in the positive direction and 58-58 to 78.6-78.6 mrad in the negative direction, designated U and V, respectively. The resulting sequence of superlayers is A, U, A, V, A, U, A, V, A.
@@
For simulation, reconstruction, calibration, and physics analysis within Belle II, a central, modular C++/Python framework, called Belle II Analysis Software Framework (basf2), is used [24, 11]. For the track reconstruction, Belle II employs a multi-step scheme implemented within basf2. The process begins with a CDC-only track-finding based on the Legendre algorithm [7], followed by attachment of SVD clusters with a combinatorial Kalman filter (CKF) [26, 27, 17, 1]. The remaining SVD hits are used to reconstruct tracks with a standalone SVD track-finder using Sector Map and cellular automaton algorithms [19, 20, 3, 18]. The results are combined and extrapolated to the PXD with a second CKF. Prior to the extrapolation between different detectors and for the final combination, the tracks are fitted using the deterministic annealing filter (DAF) provided by the genfit2 package [13, 23]. More detailed information about the Belle II track-finding algorithm is provided in Ref. [12].
@@
The benchmarking of the Kink Finder is carried out using $\tau^{+}\tau^{-}$ events, generated with KKMC [22] and TAUOLA [21] packages, and $B\bar{B}$ events, generated with PYTHIA8 [30] and EvtGen [25] packages. The full detector geometry, interactions of final state particles with the detector material, and the detector response are simulated using the GEANT4 package [5].
@@ 3 Kinks at Belle II
The latter case is relevant to the PID fake rate, as mentioned in Sec. 1. In around 10% of in-flight decay events of kaons and pions from $\tau^{-}\rightarrow K^{-}\nu_{\tau}$ and $\tau^{-}\rightarrow\pi^{-}\nu_{\tau}$ decays, the daughter track cannot be reconstructed since it leaves too few hits inside the CDC. A similar fraction of in-flight decay events is obtained for average kaon and pion in $B\bar{B}$ events. Although some of the kink processes produce more than one charged daughter particle, it is usually possible to reconstruct and combine only one of them with the mother track, e.g., it is quite rare that more than one daughter pion is reconstructed from the $K^{-}\rightarrow\pi^{-}\pi^{+}\pi^{-}$ decay. Examples of MC simulated events with kinks inside the Belle II CDC are shown in Fig. 1. The blue and red hits show the reconstructed mother and daughter tracks, respectively, while yellow hits represent all reconstructed hits inside the CDC, including the background.
@@
There are around 40% of kaon and 70% of pion events containing in-flight decays, in which there is one reconstructed track arising from a combination of mother and daughter hits. Often, the hits attributed to either the mother or the daughter dominate in such combined track, allowing for its relation to the dominant particle at the MC level. Therefore, some of the combined tracks are included in the calculation of the reconstruction efficiencies above. Around 40% of the combined tracks can neither be attributed to the mother nor to the daughter. This is because the ratio of the hit contributions is almost 50%/50%, whereas the Belle II definition requires at least 66% of hits from a single particle to create a relation with it [12].
@@ 4 Kink Finder algorithm
4.1 Reconstructed track pairs
First, we describe the case with both mother and daughter tracks reconstructed individually. To optimize the execution time of the algorithm, we perform a preselection of tracks that are candidates to form a kink. The mother track candidate should start in the vicinity of the IP and, if it has CDC hits, end inside the CDC. The daughter track candidate should start (or end, if the assigned charge of the track is wrong) inside the CDC and its point of closest approach to the IP should be at a minimum distance.
@@ 4.1 Reconstructed track pairs
    The starting point of the daughter track is close to the endpoint of the mother track in 3D [see Fig. 1(a)].
@@
    The endpoint of the daughter track is close to the endpoint of the mother track in 3D [see Fig. 1(b)]. This configuration covers events with wrong charge assignment of the daughter track.
@@
    The extrapolated daughter track is close to the endpoint of the mother track in 3D [see Fig. 1(c)]. This configuration covers events in which the daughter track intermediate hits are lost.
@@
Several assisting algorithms are applied to improve the kink parameters resolution in the case of two separately reconstructed tracks. One example is hit reassignment between tracks and daughter track refitting with better initial values (seeds). We start with track pairs selected by the first three configurations (described above), all of which have good $z$-coordinate resolution for the daughter track. For these pairs, the processing begins with a vertex fit to find the location of the kink. The fitting procedure is described in Sec. 4.1.1. In the case of poor $z$-coordinate resolution of the daughter track, the track is refitted using the time and position information from the endpoint of the mother track. If the track charges are correctly reconstructed, and the start point and end point of the tracks are close to each other (normal ordering), the poor $z$-coordinate resolution of the daughter track may be due to wrong assignment of the first stereo superlayer in the CDC. We then try to refit the daughter track without these hits. Using the new fit result for the daughter, the vertex fit is performed.
@@
While the DAF fitting algorithm from the genfit2 package is used as the default track fitter at Belle II, we use the Reference Kalman Filter from the same package [13, 23] for the fit of the flipped tracks and the refit of the daughter tracks with missing superlayer hits, as it shows a better performance.
@@
      The p-value of the fitted combined track [12] is greater than the initial mother track’s p-value.
@@ 4.1.1 Vertex fit
The fit of the vertex position is performed using a kinematic fitter algorithm implemented in the Belle II Software. It is based on the least squares method and described in Ref. [31], Appendix D. The genfit2 package also provides an algorithm for the vertex fitting, GFRave [13, 23, 32], which is applied at Belle II for the V0 ($K^{0}_{S}$, $\Lambda$, and externally converted $\gamma$) vertex reconstruction. However, we do not use it in the Kink Finder since it does not have an option for setting an initial value. Without such an option, the vertex fit fails for a noticeable amount of kinks.
@@ 4.2 Combined tracks
  1. 1.
    The p-value of the track fit [12] must be lower than the threshold, which is set several orders of magnitude larger than the loose threshold used in general track fitting. The value used in the Kink Finder is optimized for the experimental conditions to suppress the splitting of one true track into two false tracks.
  2. 2.
  3. 3.
Then, the combined track candidates are split using the algorithm that creates a pair of mother and daughter tracks using the hits of the initial track.
@@ 5 Performance
The Kink Finder performance for in-flight decays is studied using MC samples of the $\tau^{-}\rightarrow K^{-}\nu_{\tau}$ and $\tau^{-}\rightarrow\pi^{-}\nu_{\tau}$ events with a total number of $8\times 10^{4}$ kaon and $13\times 10^{4}$ pion in-flight decays inside the CDC. The performance for general cases of kinks is studied using MC samples of generic $\tau$-pair and $B\bar{B}$ decays with $10^{5}$ events in each sample. In both studies, we account for beam-induced background, which consists of spurious detector hits and particles produced in secondary interactions with the detector and support material outside the IP region [28]. The level of this background strongly depends on the SuperKEKB luminosity and its conditions. In our studies, we use a configuration that is considered typical for current Belle II operations. Although the Kink Finder has been included in the Belle II tracking software, it has not been applied for data processing yet, and therefore, there is no performance study using data available at the time of writing.